OpenStudy (anonymous):

3x=5(mod 7); find x?

Nnesha (nnesha):

@ganeshie8 mod^

OpenStudy (amistre64):

3x = 5 (mod 7) this is defined as: 3x = 7k + 5, for any integer k x = (7k+5)/3; let k= 0,1,2,3,4,... 5/3 12/3 = 4 mod 7 19/3 26/3 33/3 = 11 = 4 mod 7 .... seems to be a pattern

OpenStudy (amistre64):

another way to see it 3x = 5 (mod 7) 3x-5 = 5-5 (mod 7) 3x-5 = 0 (mod 7) means that 3x-5 is a multiple of 7 ---------------------------- 3x+2 = 5+2 (mod 7) works just as well i spose